# Unitarization of the Sommerfeld enhancement through the renormalization group

### v1

@@
When a pair of dark matter particles interacts via a long-range force mediated by a light particle, their nonrelativistic annihilation cross section can be significantly enhanced—a phenomenon known as the Sommerfeld enhancement. This enhancement exhibits resonant behavior if the long-range potential supports shallow bound states or narrow resonances, which can lead to violations of the partial-wave unitarity bound. We identify the origin of this pathological behavior as the emergence of secular terms in perturbative expansions associated with low-energy composite states of the long-range potential. To address this issue, we propose a renormalization group improvement of the perturbative series. The resulting improved amplitude provides a unitarity-consistent form of the Sommerfeld enhancement, with its poles acquiring a imaginary part that reflects the decay width of the annihilating bound states. We also briefly discuss the implications of our approach from the perspective of Wilsonian renormalization group, and comment on its potential application to higher-order annihilation processes such as bound-state formation.
@@ 1 Introduction
Sommerfeld enhancement exhibits a resonant behavior and becomes significantly amplified when the parameters of the long-range force approach certain critical values. This occurs when the spectrum generated by the long-range interaction contains shallow bound states or narrow  [5, 7, 8]. However, such an excessive amplification of the annihilation cross section is, in fact, inconsistent with quantum theory. Specifically, it violates the unitarity of quantum mechanics—more precisely, it leads to a contradiction with the unitarity bound on the annihilation cross section, which is a direct consequence of unitarity [5, 7, 9]. For example, in the case of $s$-wave annihilation, if there exists a shallow bound state, the annihilation cross section including the Sommerfeld enhancement scales as $v^{-2}$ with respect to the dark matter velocity $v$, whereas the unitarity bound scales as $v^{-1}$, leading to an inconsistency in the small-$v$ regime. Accepting the excessively enhanced Sommerfeld effect without modification may result in a misrepresentation of the model’s validity. Accordingly, it is of crucial importance to unitarize the Sommerfeld enhancement—namely, to reformulate it in a manner consistent with the unitarity.
@@
Singular perturbation theory is a well-known issue, and it is established that it can be systematically addressed using renormalization group methods [16, 17]. In this paper, we propose a unitarization scheme for the Sommerfeld enhancement, in which we resum the secular terms that appear in the perturbative expansion of the self-scattering amplitude using the renormalization group approach. The ingredients required for this procedure are the short-distance scattering amplitude from the UV theory and the phase shift induced by the long-range interaction. The resulting improved scattering amplitude, as well as the modified Sommerfeld factor derived from it, are both manifestly consistent with unitarity. Furthermore by analyzing its analytic structure, we confirm the appearance of an imaginary part in the bound-state energy—i.e., a decay width—thus realizing the physical interpretation proposed in Ref.[5] in a concrete manner.
The organization of this paper is as follows. In the next section (Section 2), we first introduce the basic tools of nonrelativistic scattering theory, which are then used to derive the conventional Sommerfeld enhancement. Through this derivation, the origin of the unitarity violation becomes manifest. In the following Section 3, motivated by this breakdown of unitarity, we introduce a renormalization group approach. We then analyze the properties of the resulting improved scattering amplitude, confirming that it is consistent with unitarity and that the positions of its poles are shifted to include the decay widths of the corresponding bound states. Subsequently in Section 4, we employ insights from Wilsonian renormalization group theory to further discuss our results. We also comment on the approach based on directly solving the Schrödinger equation found in [9, 10]. A brief discussion is also given on bound-state formation [18, 19, 20] as a higher-order annihilation process. In Section 5 is devoted to the conclusion. Appendix  A provides a summary of the Bessel functions used in the nonrelativistic scattering theory, and Appendix B collects technical details omitted in the main text.
@@ 2.1 Basic concepts in scattering theory
We begin with the most fundamental quantity in scattering theory: the $S$-matrix. Consider a pair of distinguishable quantum mechanical particles with reduced mass $\mu$. [^2] In the case of identical particles, additional numerical factors (such as 2) may appear in certain expressions [11, 10]. These particles evolve under the free Hamiltonian $H_{0}$ in the distant past, interact via a interaction $V$ localized around time $t=0$, and then continue to evolve freely under $H_{0}$ in the distant future. The $S$-matrix element for the transition from an asymptotically free state to another, labeled by $\alpha$ and $\beta$, is defined as
@@
where $\delta_{\ell}(p)$ is the phase shift for the angular momentum $\ell$ and is a real function due to the unitarity of the $S$-matrix. Here, we parametrize the energy $E$ using the magnitude of the momentum $p=2\mu E$. The partial-wave decomposition of the scattering amplitude is then given by
@@
One of the major utilities of the analyticity of the amplitude lies in the statement that its singularities are believed to correspond to spectral properties of the system. Specifically, bound states correspond to poles of the scattering amplitude. Moreover, unitarity, or equivalently the optical theorem, suggests that the scattering amplitude possesses branch cuts with thresholds as branch points. Indeed, requiring Hermitian analyticity of the amplitude, $f_{\ell}(E^{*})=[f_{\ell}(E)]^{*}$ leads, via the optical theorem, to the presence of an imaginary part in the scattering region, which in turn results in discontinuities and hence the presence of branch cuts. In the next section, we will see explicitly how, in non-relativistic quantum mechanics, the $S$-matrix or scattering amplitude can be analytically continued into complex energy or complex momentum, realize these structures.
@@ 2.2 The Jost function and the analytic properties of the $S$-matrix
The solution of (2.10) we need for the scattering problem is completely determined by its boundary conditions and they are, at the origin, the regularity of the wave function $u_{\ell,p}(0)=0$  [^5] The irregular solution implies that the Hamiltonian includes a delta-function potential localized at the origin [21]. Imposing the regularity condition excludes the existence of such a singular operator. Conversely, for a Schrödinger equation with a singular potential, the solution necessarily contains a irregular component.and, at infinity, that the wave function is a superposition of an incoming wave and an outgoing wave,
$u_{\ell,p}(p)\to\frac{i}{2}[h_{\ell}^{-}(pr)-s_{\ell}(p)h^{+}_{\ell}(pr)]\quad(r\to\infty),$ (2.11)
where $h_{\ell}^{\mp}(pr)$ are the Riccati-Hankel functions, which are the solutions of the Scrödinger equation (2.10) with $V(r)=0$. Their asymptotic behavior at infinity are given by $h_{\ell}^{\mp}(pr)\sim e^{\mp ipr}$, that is, $h^{-}_{\ell}(pr)$ and $h^{+}_{\ell}(pr)$ correspond to the in-coming and outgoing waves, respectively (See Appendix A). Therefore, in the formulation of the scattering problem based on the Schrödinger equation, the $S$-matrix is extracted from the relative weight of the incoming and outgoing waves.
@@
Since both the regular solution $\phi_{\ell,p}(r)$ and the scattering wave function $u_{\ell,p}(r)$ are regular at the origin, they must be proportional to each other in the physical region $p>0$, where the scattering solution $u_{\ell,p}(r)$ exists, and we can find that those are related by $u_{\ell,p}(r)=\phi_{\ell,p}(r)/\mathscr{J}_{\ell}(p)$ by compering their asymptotic forms (2.11) and (2.13). Therefore, the above definition agrees with the original $S$-matrix on the positive real axis and provides its natural analytic continuation into the complex $p$ plane. In addition, using the relation $\mathscr{J}_{\ell}(p)=[\mathscr{J}_{\ell}(-p^{*})]^{*}$, which follows from (2.14) (see Appendix B for details), we obtain $s_{\ell}(p)=e^{2i\delta_{\ell}(p)}=\mathscr{J}_{\ell}^{*}(p)/\mathscr{J}_{\ell}(p)$, indicating that the phase of the Jost function in the physical region $p>0$ is given by $\delta_{\ell}(p)$. [^8] There is an ambiguity of $2\pi i$ in the phase shift $\delta_{\ell}(p)$, and here we choose it to be $\delta_{\ell}(\infty)=0$. However, this choice does not affect the subsequent discussion.
It is noteworthy to consider the case where the analytically continued $S$-matrix or scattering amplitude (2.15) has poles in the complex $p$-plane. Suppose that the $S$-matrix $s_{\ell}(p)=\mathscr{J}_{\ell}^{*}(p)/\mathscr{J}_{\ell}(p)$ has a pole on the positive imaginary axis of the $p$-plane; that is, the Jost function vanishes at some $p=i\kappa,\kappa>0$. In such a situation, as can be seen in equation (2.13), the incoming wave component of the regular solution $h_{\ell}^{-}(pr)$ vanishes at infinity, while the outgoing wave component remains in decaying form: $\phi_{\ell,p}(r)\sim e^{-\kappa r}$. This is precisely the boundary condition imposed on the Schrödinger equation (2.10) when solving for bound states. Conversely, for any value of $p$ where a bound-state wave function exists, it must be proportional to the regular solution due to regularity at the origin, and thus the growing incoming wave component at infinity must be absent. Therefore, a zero of the Jost function on the positive imaginary axis, that is, a pole of the $S$-matrix or the scattering amplitude [^9] The zeros of the Jost function do not correspond one-to-one with the poles of the $S$-matrix. This is because some poles may arise from divergences in the numerator, and such poles are referred to as redundant poles. In fact, the $S$-matrix for an exponential potential includes redundant poles., corresponds to a bound state, and the associated energy $E_{=}-\kappa^{2}/2\mu$ gives the binding energy of that state.
@@
Figure 1: Analytic structure of the scattering amplitude on the complex $p$ and $E=\sqrt{2\mu p}$ plane. The left panel shows the complex $p$-plane, while the middle and right panels represent the physical and unphysical sheets of the energy Riemann surface, respectively. The two sheets are connected across the branch cut, depicted by the wavy lines along the positive real axis. Sample paths illustrating analytic continuation around the branch point, as well as the corresponding trajectories in the $p$-plane and possible pole singularities, are also indicated. See the main text for details.
The analytic structure of the scattering amplitude and the Jost function is naturally described on both the complex $p$-plane and the associated Riemann surface of the energy $E=p^{2}/2\mu$, a representative example of which is illustrated in Fig. 1, where the upper (lower) half of the complex $p$-plane corresponds to the first (second) Riemann sheet of the energy, denoted $E({\rm I})$ and $E({\rm II})$. In order to analytically connect these sheets, a branch cut —usually referred to as the right-hand cut or the unitarity cut — is placed along the positive real axis of the complex $E$-plane, consistent with the requirement that scattering occurs for $p>0$ and with the constraint imposed by the optical theorem. [^10] Depending on the potential, the scattering amplitude may exhibit a left-hand cut, known as a dynamical cut, in contrast to the right-hand cut required by unitarity. For example, the Yukawa potential $V(r)=\alpha e^{-mr}/\mu$ generates such a cut starting from $E_{L}=-m^{2}/8\mu$. In relativistic scattering, due to crossing symmetry, the unitarity cut automatically implies the existence of the dynamical cut. A sample path encircling the branch point at $E=0$ is also shown in Fig. 1. The path starts on the upper edge of the branch cut on $E({\rm I})$, moves to the lower edge via the red arc, continues to the upper edge of $E({\rm II})$, and then to its lower edge along the blue dotted arc, finally returning to the starting point on $E({\rm I})$. The corresponding loop is also illustrated on the complex $p$-plane. The physical scattering amplitude is then identified with the boundary value on the upper (lower) edge of the branch cut on $E({\rm I})$ ($E({\rm II})$), and the two sheets are thus referred to as the physical and unphysical sheets, respectively.
@@ 2.3 Sommerfeld enhancement
where $V_{L}(r)$ is a long-range component representing their persistent interaction, and $V_{S}(r)$ is a short-range component responsible for inducing their annihilation. Schematically, the Sommerfeld effect arises due to the distortion of the plane wave associated with the annihilation via $V_{S}(r)$ caused by the long-range interaction $V_{L}(r)$. As a concrete example of those potentials encountered in actual problems, the long-range part is typically modeled by a Yukawa interaction, while the short-range part is assumed to consist of a delta function and its higher derivatives. For instance, in a system where a particle $\varphi$ of mass $M$ couples to a light vector mediator $A$ of mass $m$ with the structure constant $\alpha$, the dynamics of the interpolating field for the two-body state $\Psi(x,\vec{r})\sim\varphi^{\dagger}(t,\vec{x}+\vec{r}/2)\varphi(t,\vec{x}-\vec{r}/2)$—after integrating out the hard scale $M$ and the soft scale $Mv$ with $v$ being the relative velocity—is described at the leading order by the following potential non-relativistic Lagrangian [30, 31, 32]:
$\mathcal{L}_{\rm pNR}\simeq\int d^{3}r\Psi^{\dagger}(x,\vec{r})\left(i\partial_{t}+\frac{\nabla^{2}_{\vec{x}}}{4M}+\frac{\nabla_{r}^{2}}{m}+\alpha\frac{e^{-mr}}{r}+i\frac{2\pi\alpha^{2}}{M^{2}}\delta(\vec{r})\right)\Psi(x,\vec{r})+\cdots.$ (2.20)
From the Lagrangian (2.20), one can move to the Hamiltonian formalism and, by projecting the Fock space spanned by two-body states onto the Hilbert space consisting of single two-body state, obtain a Hamiltonian with the potentials $V_{L}(r)=-\alpha e^{-mr}/r$, $V_{S}(r)=-i2\pi\alpha^{2}\delta(\vec{r})/M^{2}$. The long-range and short-range potentials originate from the contributions of the soft and hard scales, respectively.
Let us describe the Sommerfeld effect as a scattering problem under the potential (2.19). First, the Born expansion, which treats the potential as a mere perturbation, fails to yield a reliable result. In fact, under the first Born approximation with the Yukawa potential $V_{L}(r)=-\alpha^{-mr}/r$, the scattering amplitude behaves as $f(\vec{q})=-2\alpha m/(\vec{q}^{2}+m^{2})$ for the momentum transfer $\vec{q}$, and becomes singular in the non-relativistic limit $\vec{q}\to 0$ when the mass of the mediator $m$ is small. Moreover, due to the unitarity requirement, higher-order Born terms also exhibit similar singular behavior. Therefore, the Born approximation is not suitable in nonrelativistic regimes, and this is precisely the origin of the Sommerfeld effect.
@@
Another possibility is the existence of a resonance state in the spectrum, namely when the Jost function has a zero in the region ${\rm Re}\,p>0$[8]. In this case, if the zero lies not too far from the real axis, that is, the decay width of the resonance is sufficiently small, one expects the Sommerfeld factor to exhibit a resonance feature, much like how the elastic scattering cross section displays a Breit–Wigner–type resonance. However, we must keep in mind that these sharp peaks may be misleading, and the actual annihilation cross section may not exhibit such behavior. This is because, in deriving Eq. (2.25), we have neglected the $\mathcal{O}(\mathscr{J}_{\ell}(p)^{-4})$ term associated with the square of the UV amplitude, which is not a justified approximation in the momentum region where such peaks appear. This observation naturally leads to the subject of the next section, which concerns the unitarization of the Sommerfeld effect.

### v2

@@
When a pair of dark matter particles interacts via a long-range force mediated by a light particle, their nonrelativistic annihilation cross section can be significantly enhanced—a phenomenon known as the Sommerfeld enhancement. This enhancement exhibits resonant behavior if the long-range potential supports shallow bound states or narrow resonances, which can lead to violations of the partial-wave unitarity bound. We identify the origin of this pathological behavior as the emergence of secular terms in perturbative expansions associated with low-energy composite states of the long-range potential. To address this issue, we propose a renormalization group improvement of the perturbative series. The resulting improved amplitude provides a unitarity-consistent form of the Sommerfeld enhancement, with its poles acquiring an imaginary part that reflects the decay width of the annihilating bound states. We also briefly discuss the implications of our approach from the perspective of Wilsonian renormalization group, and comment on its potential application to higher-order annihilation processes such as bound-state formation.
@@ 1 Introduction
Sommerfeld enhancement exhibits a resonant behavior and becomes significantly amplified when the parameters of the long-range force approach certain critical values. This occurs when the spectrum generated by the long-range interaction contains shallow bound states or narrow resonances [5, 7, 8]. However, such an excessive amplification of the annihilation cross section is, in fact, inconsistent with quantum theory. Specifically, it violates the unitarity of quantum mechanics—more precisely, it leads to a contradiction with the unitarity bound on the annihilation cross section, which is a direct consequence of unitarity [5, 7, 9]. For example, in the case of $s$-wave annihilation, if there exists a shallow bound state, the annihilation cross section including the Sommerfeld enhancement scales as $v^{-2}$ with respect to the dark matter velocity $v$, whereas the unitarity bound scales as $v^{-1}$, leading to an inconsistency in the small-$v$ regime. Accepting the excessively enhanced Sommerfeld effect without modification may result in a misrepresentation of the model’s validity. Accordingly, it is of crucial importance to unitarize the Sommerfeld enhancement—namely, to reformulate it in a manner consistent with unitarity.
@@
Singular perturbation theory is a well-known issue, and it is established that it can be systematically addressed using renormalization group methods [16, 17]. In this paper, we propose a unitarization scheme for the Sommerfeld enhancement, in which we resum the secular terms that appear in the perturbative expansion of the self-scattering amplitude using the renormalization group approach. The ingredients required for this procedure are the short-distance scattering amplitude from the UV theory and the phase shift induced by the long-range interaction. The resulting improved scattering amplitude, as well as the modified Sommerfeld factor derived from it, are both manifestly consistent with unitarity. Furthermore, by analyzing its analytic structure, we confirm the appearance of an imaginary part in the bound-state energy—i.e., a decay width—thus realizing the physical interpretation proposed in Ref. [5] in a concrete manner.
The organization of this paper is as follows. In the next section (Section 2), we first introduce the basic tools of nonrelativistic scattering theory, which are then used to derive the conventional Sommerfeld enhancement. Through this derivation, the origin of the unitarity violation becomes manifest. In the following Section 3, motivated by this breakdown of unitarity, we introduce a renormalization group approach. We then analyze the properties of the resulting improved scattering amplitude, confirming that it is consistent with unitarity and that the positions of its poles are shifted to include the decay widths of the corresponding bound states. Subsequently, in Section 4, we employ insights from Wilsonian renormalization group theory to further discuss our results. We also comment on the approach based on directly solving the Schrödinger equation found in [9, 10]. A brief discussion is also given on bound-state formation [18, 19, 20] as a higher-order annihilation process. Section 5 is devoted to the conclusion. Appendix  A provides a summary of the Bessel functions used in the nonrelativistic scattering theory, and Appendix B collects technical details omitted in the main text. Appendix C numerically confirms that our method reproduces the results of [9, 10].
@@ 2.1 Basic concepts in scattering theory
We begin with the most fundamental quantity in scattering theory: the $S$-matrix. Consider a pair of distinguishable quantum mechanical particles with reduced mass $\mu$. [^2] In the case of identical particles, additional numerical factors (such as 2) may appear in certain expressions [11, 10]. These particles evolve under the free Hamiltonian $H_{0}$ in the distant past, interact via an interaction $V$ localized around time $t=0$, and then continue to evolve freely under $H_{0}$ in the distant future. The $S$-matrix element for the transition from an asymptotically free state to another, labeled by $\alpha$ and $\beta$, is defined as
@@
where $\delta_{\ell}(p)$ is the phase shift for the angular momentum $\ell$ and is a real function due to the unitarity of the $S$-matrix. Here, we parametrize the energy $E$ using the magnitude of the momentum $E=p^{2}/2\mu$. The partial-wave decomposition of the scattering amplitude is then given by
@@
One of the major utilities of the analyticity of the amplitude lies in the statement that its singularities are believed to correspond to spectral properties of the system. Specifically, bound states correspond to poles of the scattering amplitude. Moreover, unitarity, or equivalently the optical theorem, suggests that the scattering amplitude possesses branch cuts with thresholds as branch points. Indeed, requiring Hermitian analyticity of the amplitude, $f_{\ell}(E^{*})=[f_{\ell}(E)]^{*}$ leads, via the optical theorem, to the presence of an imaginary part in the scattering region, which in turn results in discontinuities and hence the presence of branch cuts. In the next section, we will see explicitly how, in non-relativistic quantum mechanics, the $S$-matrix or scattering amplitude can be analytically continued into complex energy or complex momentum, and realize these structures.
@@ 2.2 The Jost function and the analytic properties of the $S$-matrix
The solution of (2.10) we need for the scattering problem is completely determined by its boundary conditions and they are, at the origin, the regularity of the wave function $u_{\ell,p}(0)=0$  [^5] The irregular solution implies that the Hamiltonian includes a delta-function potential localized at the origin [21]. Imposing the regularity condition excludes the existence of such a singular operator. Conversely, for a Schrödinger equation with a singular potential, the solution necessarily contains an irregular component.and, at infinity, that the wave function is a superposition of an incoming wave and an outgoing wave,
$u_{\ell,p}(r)\to\frac{i}{2}[h_{\ell}^{-}(pr)-s_{\ell}(p)h^{+}_{\ell}(pr)]\quad(r\to\infty),$ (2.11)
where $h_{\ell}^{\mp}(pr)$ are the Riccati-Hankel functions, which are the solutions of the Schrödinger equation (2.10) with $V(r)=0$. Their asymptotic behavior at infinity are given by $h_{\ell}^{\mp}(pr)\sim e^{\mp ipr}$, that is, $h^{-}_{\ell}(pr)$ and $h^{+}_{\ell}(pr)$ correspond to the incoming and outgoing waves, respectively (See Appendix A). Therefore, in the formulation of the scattering problem based on the Schrödinger equation, the $S$-matrix is extracted from the relative weight of the incoming and outgoing waves.
@@
Since both the regular solution $\phi_{\ell,p}(r)$ and the scattering wave function $u_{\ell,p}(r)$ are regular at the origin, they must be proportional to each other in the physical region $p>0$, where the scattering solution $u_{\ell,p}(r)$ exists, and we can find that those are related by $u_{\ell,p}(r)=\phi_{\ell,p}(r)/\mathscr{J}_{\ell}(p)$ by comparing their asymptotic forms (2.11) and (2.13). Therefore, the above definition agrees with the original $S$-matrix on the positive real axis and provides its natural analytic continuation into the complex $p$ plane. In addition, using the relation $\mathscr{J}_{\ell}(p)=[\mathscr{J}_{\ell}(-p^{*})]^{*}$, which follows from (2.14) (see Appendix B for details), we obtain $s_{\ell}(p)=e^{2i\delta_{\ell}(p)}=\mathscr{J}_{\ell}^{*}(p)/\mathscr{J}_{\ell}(p)$, indicating that the phase of the Jost function in the physical region $p>0$ is given by $\delta_{\ell}(p)$. [^8] There is an ambiguity of $2\pi i$ in the phase shift $\delta_{\ell}(p)$, and here we choose it to be $\delta_{\ell}(\infty)=0$. However, this choice does not affect the subsequent discussion.
It is noteworthy to consider the case where the analytically continued $S$-matrix or scattering amplitude (2.15) has poles in the complex $p$-plane. Suppose that the $S$-matrix $s_{\ell}(p)=\mathscr{J}_{\ell}^{*}(p)/\mathscr{J}_{\ell}(p)$ has a pole on the positive imaginary axis of the $p$-plane; that is, the Jost function vanishes at some $p=i\kappa,\kappa>0$. In such a situation, as can be seen in equation (2.13), the incoming wave component of the regular solution $h_{\ell}^{-}(pr)$ vanishes at infinity, while the outgoing wave component remains in decaying form: $\phi_{\ell,p}(r)\sim e^{-\kappa r}$. This is precisely the boundary condition imposed on the Schrödinger equation (2.10) when solving for bound states. Conversely, for any value of $p$ where a bound-state wave function exists, it must be proportional to the regular solution due to regularity at the origin, and thus the growing incoming wave component at infinity must be absent. Therefore, a zero of the Jost function on the positive imaginary axis, that is, a pole of the $S$-matrix or the scattering amplitude [^9] The zeros of the Jost function do not correspond one-to-one with the poles of the $S$-matrix. This is because some poles may arise from divergences in the numerator, and such poles are referred to as redundant poles. In fact, the $S$-matrix for an exponential potential includes redundant poles., corresponds to a bound state, and the associated energy $E=-\kappa^{2}/2\mu$ gives the binding energy of that state.
@@
Figure 1: Analytic structure of the scattering amplitude on the complex $p$ and $E=p^{2}/2\mu$ plane. The left panel shows the complex $p$-plane, while the middle and right panels represent the physical and unphysical sheets of the energy Riemann surface, respectively. The two sheets are connected across the branch cut, depicted by the wavy lines along the positive real axis. Sample paths illustrating analytic continuation around the branch point, as well as the corresponding trajectories in the $p$-plane and possible pole singularities, are also indicated. See the main text for details.
The analytic structure of the scattering amplitude and the Jost function is naturally described on both the complex $p$-plane and the associated Riemann surface of the energy $E=p^{2}/2\mu$, a representative example of which is illustrated in Fig. 1, where the upper (lower) half of the complex $p$-plane corresponds to the first (second) Riemann sheet of the energy, denoted $E({\rm I})$ and $E({\rm II})$. In order to analytically connect these sheets, a branch cut —usually referred to as the right-hand cut or the unitarity cut — is placed along the positive real axis of the complex $E$-plane, consistent with the requirement that scattering occurs for $p>0$ and with the constraint imposed by the optical theorem. [^10] Depending on the potential, the scattering amplitude may exhibit a left-hand cut, known as a dynamical cut, in contrast to the right-hand cut required by unitarity. For example, the Yukawa potential $V(r)=\alpha e^{-m_{V}r}/r$ generates such a cut starting from $E_{L}=-m_{V}^{2}/8\mu$. In relativistic scattering, due to crossing symmetry, the unitarity cut automatically implies the existence of the dynamical cut. A sample path encircling the branch point at $E=0$ is also shown in Fig. 1. The path starts on the upper edge of the branch cut on $E({\rm I})$, moves to the lower edge via the red arc, continues to the upper edge of $E({\rm II})$, and then to its lower edge along the blue dotted arc, finally returning to the starting point on $E({\rm I})$. The corresponding loop is also illustrated on the complex $p$-plane. The physical scattering amplitude is then identified with the boundary value on the upper (lower) edge of the branch cut on $E({\rm I})$ ($E({\rm II})$), and the two sheets are thus referred to as the physical and unphysical sheets, respectively.
@@ 2.3 Sommerfeld enhancement
where $V_{L}(r)$ is a long-range component representing their persistent interaction, and $V_{S}(r)$ is a short-range component responsible for inducing their annihilation. Schematically, the Sommerfeld effect arises due to the distortion of the plane wave associated with the annihilation via $V_{S}(r)$ caused by the long-range interaction $V_{L}(r)$. As a concrete example of those potentials encountered in actual problems, the long-range part is typically modeled by a Yukawa interaction, while the short-range part is assumed to consist of a delta function and its higher derivatives. For instance, in a system where a particle $\varphi$ of mass $M$ couples to a light vector mediator $A$ of mass $m_{V}$ with the structure constant $\alpha$, the dynamics of the interpolating field for the two-body state $\Psi(x,\vec{r})\sim\varphi^{\dagger}(t,\vec{x}+\vec{r}/2)\varphi(t,\vec{x}-\vec{r}/2)$—after integrating out the hard scale $M$ and the soft scale $Mv$ with $v$ being the relative velocity—is described at the leading order by the following potential non-relativistic Lagrangian [30, 31, 32, 33]:
$\mathcal{L}_{\rm pNR}\simeq\int d^{3}r\Psi^{\dagger}(x,\vec{r})\left(i\partial_{t}+\frac{\nabla^{2}_{\vec{x}}}{4M}+\frac{\nabla_{r}^{2}}{M}+\alpha\frac{e^{-m_{V}r}}{r}+i\frac{2\pi\alpha^{2}}{M^{2}}\delta(\vec{r})\right)\Psi(x,\vec{r})+\cdots.$ (2.20)
From the Lagrangian (2.20), one can move to the Hamiltonian formalism and, by projecting the Fock space spanned by two-body states onto the Hilbert space consisting of single two-body state, obtain a Hamiltonian with the potentials $V_{L}(r)=-\alpha e^{-m_{V}r}/r$, $V_{S}(r)=-i2\pi\alpha^{2}\delta(\vec{r})/M^{2}$. The long-range and short-range potentials originate from the contributions of the soft and hard scales, respectively.
Let us describe the Sommerfeld effect as a scattering problem under the potential (2.19). First, the Born expansion, which treats the potential as a mere perturbation, fails to yield a reliable result. In fact, under the first Born approximation with the Yukawa potential $V_{L}(r)=-\alpha e^{-m_{V}r}/r$, the scattering amplitude behaves as $f(\vec{q})=-2\alpha m/(\vec{q}^{2}+m^{2})$ for the momentum transfer $\vec{q}$, and becomes singular in the non-relativistic limit $\vec{q}\to 0$ when the mass of the mediator $m$ is small. Moreover, due to the unitarity requirement, higher-order Born terms also exhibit similar singular behavior. Therefore, the Born approximation is not suitable in nonrelativistic regimes, and this is precisely the origin of the Sommerfeld effect.
@@
Another possibility is the existence of a resonance state in the spectrum, namely when the Jost function has a zero in the region ${\rm Re}\,p>0$[8]. In this case, if the zero lies not too far from the real axis, that is, the decay width of the resonance is sufficiently small, one expects the Sommerfeld factor to exhibit a resonance feature when the energy of scattering particles approaches the resonance energy, much like how the elastic scattering cross section displays a Breit–Wigner–type resonance. However, we must keep in mind that the actual annihilation cross section as a function of momentum $p$ may not exhibit such resonance behavior. This is because, in deriving Eq. (2.25), we have neglected the $\mathcal{O}(\mathscr{J}_{\ell}(p)^{-4})$ term arising from the square of the UV amplitude, and this approximation is no longer valid in the momentum region close to the corresponding resonance energy. This observation naturally leads to the subject of the next section, which concerns the unitarization of the Sommerfeld effect.